if a=sec theta - tan theta, b=cosec theta + cot theta; calculate a in terms of b

Grade:11

3 Answers

Sayam
6 years ago
b=cosec \Theta +cot\Theta=(1+cos\Theta) / sin\Theta= cot\Theta/2
a=sec \Theta-tan \Theta=(1-sin \Theta)/cos \Theta
=1-cos(\pi/2- \Theta)/sin(\pi/2- \Theta)
=2sin2(\pi/4- \Theta/2) / 2sin(\pi/4- \Theta/2) cos(\pi/4- \Theta/2)
=tan(\pi/4- \Theta/2)=(1-tan \Theta/2)  /  (1+tan\Theta/2)
=(1-1/b) / (1+1/b)
=b-1/b+1......  ans
hope this clears your doubt
Dhjdjddj
3 years ago
Hhshshajjajaksjsjdhdhdhdhhdhehehdjdjskskwowiwhhsgsgehshdjdbdbdussksauwowanvshewksgeieheheeidhwhshdhdhdhdhdududuhdbdbdbdbdbdhhshdhdbdhshudhdhdvsgsyhshsgshshsbbshsuajwosodebeuehejodkdkdjehdjskhdhehdhhdhehdhdhdkiaoqqpqlernnsozxhyxs duelsshsobsekendjssbbwgwwhdkdhd ho ldvdkeveieeoeur elabqkeeueled
Vikas TU
3 years ago
Dear student 
Please do not misuse the forum ,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,